OpenStudy (anonymous):

The surface areas of two similar solids are 318 yd2 and 1004 yd2. The volume of the larger solid is 1730 yd3. What is the volume of the smaller solid? A) 308 yd3 B) 822 yd3 C) 974 yd3 D) 1,730 yd3 Thanks!

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

I'm having trouble mastering your instructions...

OpenStudy (anonymous):

1) take the ratio 2) take the square root 3) cube it that is your new ratio

OpenStudy (anonymous):

thanks man!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i can check if you like

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

17.83,

OpenStudy (anonymous):

31.68,

OpenStudy (anonymous):

how do i find ratio?

OpenStudy (anonymous):

\[\frac{318}{1008}=\frac{159}{502}\] \[\sqrt{\frac{159}{502}}=.52679\] rounded

OpenStudy (anonymous):

oh i went for the square roots first...

OpenStudy (anonymous):

\[.592679^3=.17825\] so did i

OpenStudy (anonymous):

so that is your ratio

OpenStudy (anonymous):

you got it but the decimal should be in front

OpenStudy (anonymous):

now i cube it?

OpenStudy (anonymous):

when I cube it i get 0.56785?

OpenStudy (anonymous):

no i did what you did take the square root then cube i did in the same order you did

OpenStudy (anonymous):

root is .56785

OpenStudy (anonymous):

then cube to get .17825

OpenStudy (anonymous):

3.164 for the other number cubed...

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

oh wait i have confused you. i am not taking the square root or the cube of the numerator and denominator

OpenStudy (anonymous):

i am taking the square root and the cube of the entire ratio

OpenStudy (anonymous):

we start again and go step by step. initial ratio is \[\frac{318}{1004}=\frac{159}{502}\] reduced. i just reduced the ratio, that is all. now i take the square root of the hole thing

OpenStudy (anonymous):

\[\sqrt{\frac{159}{502}}=.56785\] rounded

OpenStudy (anonymous):

now i cube the result and get \[.56785^3=.17825\]

OpenStudy (anonymous):

and that is the ratio for the volumes.

OpenStudy (anonymous):

in other words that is what i would multiply your larger volume to get the smaller volume

OpenStudy (anonymous):

\[1730\times .17825=308\] rounded and that is the answer

OpenStudy (anonymous):

Okay! I see, thank you so much!

OpenStudy (anonymous):

yw